OpenStudy (anonymous):

What is the sum of the solutions of 2|x - 1| - 4 = -2?

OpenStudy (anonymous):

Will give medal.

OpenStudy (anitas.):

2(|x−1|)−4=−2 add 4 to both sides. 2(|x−1|)−4+4=−2+4 2(|x−1|)=2 then divide both sides by 2. 2(|x−1|)/2=2/2 |x−1|=1 solve Absolute Value. |x−1|=1 We know eitherx−1=1orx−1=−1 x−1=1(Possibility 1) x−1+1=1+1(Add 1 to both sides) x=2 x−1=−1(Possibility 2) x−1+1=−1+1(Add 1 to both sides) x=0 Answer: x=2 or x=0

OpenStudy (anonymous):

Thank you so much!